# CauScien: Uncovering Causality in Science

### v1

@@ 5 Conclusion
We introduced OSCAR, a scalable one-shot causal discovery framework for high-dimensional multi-label event sequences, achieving in minutes what classical methods cannot compute in days. By combining pretrained autoregressive models with parallelised CI-testing, OSCAR delivers interpretable, deployment-ready causal graphs with quantified causal indicators. This brings causal discovery closer to large-scale reasoning, making it practical for real-world sequential data.
@@ Appendix B Definitions
Spirtes et al. [35]. Given a BN $\boldsymbol{U},\mathbb{G},P>,\mathbb{G}$ is faithful to $P$ if and only if every conditional independence present in $P$ is entailed by $\mathbb{G}$ and the Markov condition holds. $P$ is faithful if and only if there exist a DAG $\mathbb{G}$ such that $\mathbb{G}$ is faithful to $P$.
@@
Tsamardinos and Aliferis [37]. In a faithful BN $\boldsymbol{U},\mathbb{G},P>$, for a set of variables $\boldsymbol{Z}\subset\boldsymbol{U}$ and label $Y\in\boldsymbol{U}$, if all other variables $X\in\{\boldsymbol{X}-\boldsymbol{Z}\}$ are independent of $Y$ conditioned on $\boldsymbol{Z}$, and any proper subset of $\boldsymbol{Z}$ do not satisfy the condition, then $\boldsymbol{Z}$ is the Markov Boundary of $Y$: $\textbf{MB}(Y)$.
@@
Variables $X$ and $Y$ are said to be conditionally independent given a variable set $\boldsymbol{Z}$, if $P(X,Y|\boldsymbol{Z})=P(X|\boldsymbol{Z})P(Y|\boldsymbol{Z})$, denoted as $X\bot Y|\boldsymbol{Z}$. Inversely, $X\not\perp Y|\boldsymbol{Z}$ denotes the conditional dependence. Using the conditional mutual information [3] to measure the independence relationship, this implies that $\text{I}(X,Y|\boldsymbol{Z})=0\Leftrightarrow X\perp Y|\boldsymbol{Z}$.
@@ Appendix C Assumptions
It allows us to remove ambiguity in causal directionality. By allowing for instantaneous rates ($t_{i}=t_{i+1})$, our method differs from Granger [10] causal discovery.
@@ E.1 Proof of Lemma 1
We assume that the data is generated by the associated causal graph $\mathbb{G}$ following the sequential BN from a multi-labelled sequence $S$. And that the faithfulness assumption holds [26], meaning that all conditional independencies in the observational data are implied by the true causal graph $\mathbb{G}$. Given that the Oracle models $\text{Tf}_{x}$ and $\text{Tf}_{y}$ are trained to perfectly approximate the true conditional distributions, for any variable $U_{i}$ in the graph, we have:
@@
The joint distribution $P_{\theta_{x},\theta_{y}}$ can then be constructed using the chain rule $P_{\theta_{x},\theta_{y}}(X_{1},\cdots,X_{i},Y_{1},\cdots,Y_{c})=\prod^{i}_{k=0}P(X_{k}|Pa(X_{k}))\prod^{c}_{l}P(Y_{l}|\text{Pa}(Y_{l})$. By the faithfulness assumption [26], if the conditional independencies hold in the data, they must also hold in the causal graph $\mathbb{G}$:
@@ E.2 Proof of Lemma 2
Base Case: $\mathbf{L=1}$: Consider the BN for step $L=1$ following the Markov assumption [26] with two nodes $X_{1},Y_{j}$. Using $\text{Tf}_{x},\text{Tf}_{y}$ as Oracle Models A4, we can express the conditional probabilities for any node $U$:
@@
Heredity: For $L=i$, we obtained $M^{i}_{j}$ with the sequential BN up to step $L=i$. Now for $L=i+1$, the sequential BN has $i+2$ nodes denoted as $\boldsymbol{U^{\prime}}=(X_{1},\cdots,X_{i},X_{i+1},Y_{j})$. Using the Oracle Models A4 and following the Markov assumption [26], we can estimates the following conditional probabilities for any nodes $U\in\boldsymbol{U^{\prime}}$:
@@ F.3 MI-MCF
MI-MCF [20] was used for comparison following the official implementation at https://github.com/malinjlu/MI-MCF we used $\alpha=0.05,L=268,k_{1}=0.7,k_{2}=0.1$.
@@ G.2 Sampling Type
We performed an ablation (Tab 3) on the effect of sampling methods to estimate the expected value over all possible context $\boldsymbol{Z}$. We used one A10 GPU on a sample of the test dataset (4000 random samples) composed of 205 labels with a batch size of 4 during inference. We tested top-k sampling with $k=\{20,35\}$ [5] with and w/o a temperature scaler of $T$ to log-probabilities $\boldsymbol{\hat{x}}$ such that
@@
And a combination of top-k and a top-nucleus sampling [13] with different probability mass $p=\{0.8,1.2\}$ and finally a permutation of token position within the context c. We fixed a dynamic threshold with z score $k=3$ and performed 10 runs. Then, we reported the average and standard deviation of each classification metric and elapsed time (sec).
@@ Appendix H Discussion on Assumptions
Inter-label Effects. By definition, the labels are explained solely by events. While simplifying multi-label causal discovery, this intrinsic assumption could be relaxed in future work by using the do operator [27] to perform interventions on common causal variables of multiple labels. For example, our current framework estimates the Markov Boundaries for each label independently. However, inter-label dependencies can exist, particularly when labels share overlapping Markov Boundaries (e.g $MB_{1}=[X_{1},X_{3}],MB_{2}=[X1,X_{2}]$. We propose to investigate a ’Phase 2’ for OSCAR, focusing on inter-label dependencies through simulated interventions. For instance, if we consider a sequence $S_{1}$ of two labels $Y_{1},Y_{2}$ with the MB above, we could perform counterfactual interventions by applying $do(X_{1}=0),do(X_{3}=0)$to $S_{1}$. Then we would observe the average change in the likelihood of $Y_{1}$ which, if it is non-zero, would indicate a dependence between $Y_{1}$ and $Y_{2}$. Wu et al. [41] points out that the assumptions of these inter-label dependencies are already anchored in the Markov Boundaries; we do the same here.
NADEs. Due to the usage of flexible NADEs, we can relax common assumptions regarding data generation processes, such as Poisson Processes or SCMs. Finally, as seen in the Ablations G.1, the effectiveness of OSCAR hinges on the capacity of $\text{Tf}_{x}$ and $\text{Tf}_{y}$ to approximate true conditional probabilities (A4) and provide Oracle CI-test. While assuming Oracle tests are common in the literature [42, 17] and necessary to recover correct causal structures, this remains a strong assumption. And it is only valid to the extent that the models are perfectly trained. Especially for multi-label classification, performance may degrade in underrepresented regions of the data distribution, as we saw during the MB length comparison.
@@ Appendix I Figures
The following is the implementation of OSCAR in PyTorch [25].

### v2

@@ 5 Conclusion
We introduced OSCAR, a scalable one-shot causal discovery framework for high-dimensional multi-label event sequences, achieving in minutes what classical methods cannot compute in days.
We pointed out limitations, notably the oracle assumption, which might not hold for rare labels. Nevertheless, by combining pretrained autoregressive Transformers with vectorized CI-testing, OSCAR delivers, for the first time, interpretable causal graphs with quantified causal indicators for thousands of different events. This brings causal discovery closer to large-scale reasoning, making it practical for real-world sequential data.
@@ Appendix B Definitions
Spirtes2001CausationPA. Given a BN $\boldsymbol{U},\mathbb{G},P>,\mathbb{G}$ is faithful to $P$ if and only if every conditional independence present in $P$ is entailed by $\mathbb{G}$ and the Markov condition holds. $P$ is faithful if and only if there exist a DAG $\mathbb{G}$ such that $\mathbb{G}$ is faithful to $P$.
@@
optimal_feature_set_cd. In a faithful BN $\boldsymbol{U},\mathbb{G},P>$, for a set of variables $\boldsymbol{Z}\subset\boldsymbol{U}$ and label $Y\in\boldsymbol{U}$, if all other variables $X\in\{\boldsymbol{X}-\boldsymbol{Z}\}$ are independent of $Y$ conditioned on $\boldsymbol{Z}$, and any proper subset of $\boldsymbol{Z}$ do not satisfy the condition, then $\boldsymbol{Z}$ is the Markov Boundary of $Y$: $\textbf{MB}(Y)$.
@@
Variables $X$ and $Y$ are said to be conditionally independent given a variable set $\boldsymbol{Z}$, if $P(X,Y|\boldsymbol{Z})=P(X|\boldsymbol{Z})P(Y|\boldsymbol{Z})$, denoted as $X\bot Y|\boldsymbol{Z}$. Inversely, $X\not\perp Y|\boldsymbol{Z}$ denotes the conditional dependence. Using the conditional mutual information [cover1999elements] to measure the independence relationship, this implies that $\text{I}(X,Y|\boldsymbol{Z})=0\Leftrightarrow X\perp Y|\boldsymbol{Z}$.
@@ Appendix C Assumptions
It allows us to remove ambiguity in causal directionality. By allowing for instantaneous rates ($t_{i}=t_{i+1})$, our method differs from granger_causality causal discovery.
@@ E.1 Proof of Lemma 1
We assume that the data is generated by the associated causal graph $\mathbb{G}$ following the sequential BN from a multi-labelled sequence $S$. And that the faithfulness assumption holds [pearl_1998_bn], meaning that all conditional independencies in the observational data are implied by the true causal graph $\mathbb{G}$. Given that the Oracle models $\text{Tf}_{x}$ and $\text{Tf}_{y}$ are trained to perfectly approximate the true conditional distributions, for any variable $U_{i}$ in the graph, we have:
@@
The joint distribution $P_{\theta_{x},\theta_{y}}$ can then be constructed using the chain rule $P_{\theta_{x},\theta_{y}}(X_{1},\cdots,X_{i},Y_{1},\cdots,Y_{c})=\prod^{i}_{k=0}P(X_{k}|Pa(X_{k}))\prod^{c}_{l}P(Y_{l}|\text{Pa}(Y_{l})$. By the faithfulness assumption [pearl_1998_bn], if the conditional independencies hold in the data, they must also hold in the causal graph $\mathbb{G}$:
@@ E.2 Proof of Lemma 2
Base Case: $\mathbf{L=1}$: Consider the BN for step $L=1$ following the Markov assumption [pearl_1998_bn] with two nodes $X_{1},Y_{j}$. Using $\text{Tf}_{x},\text{Tf}_{y}$ as Oracle Models A4, we can express the conditional probabilities for any node $U$:
@@
Heredity: For $L=i$, we obtained $M^{i}_{j}$ with the sequential BN up to step $L=i$. Now for $L=i+1$, the sequential BN has $i+2$ nodes denoted as $\boldsymbol{U^{\prime}}=(X_{1},\cdots,X_{i},X_{i+1},Y_{j})$. Using the Oracle Models A4 and following the Markov assumption [pearl_1998_bn], we can estimates the following conditional probabilities for any nodes $U\in\boldsymbol{U^{\prime}}$:
@@ F.3 MI-MCF
MI-MCF [mimcf] was used for comparison following the official implementation at https://github.com/malinjlu/MI-MCF we used $\alpha=0.05,L=268,k_{1}=0.7,k_{2}=0.1$.
@@ G.2 Sampling Type
We performed an ablation (Tab 3) on the effect of sampling methods to estimate the expected value over all possible context $\boldsymbol{Z}$. We used one A10 GPU on a sample of the test dataset (4000 random samples) composed of 205 labels with a batch size of 4 during inference. We tested top-k sampling with $k=\{20,35\}$ [fan-etal-2018-hierarchical] with and w/o a temperature scaler of $T$ to log-probabilities $\boldsymbol{\hat{x}}$ such that
@@
And a combination of top-k and a top-nucleus sampling [Holtzman2020The] with different probability mass $p=\{0.8,1.2\}$ and finally a permutation of token position within the context c. We fixed a dynamic threshold with z score $k=3$ and performed 10 runs. Then, we reported the average and standard deviation of each classification metric and elapsed time (sec).
@@ Appendix H Discussion on Assumptions
Inter-label Effects. By definition, the labels are explained solely by events. While simplifying multi-label causal discovery, this intrinsic assumption could be relaxed in future work by using the do operator [pearl_2009] to perform interventions on common causal variables of multiple labels. For example, our current framework estimates the Markov Boundaries for each label independently. However, inter-label dependencies can exist, particularly when labels share overlapping Markov Boundaries (e.g $MB_{1}=[X_{1},X_{3}],MB_{2}=[X1,X_{2}]$. We propose to investigate a ’Phase 2’ for OSCAR, focusing on inter-label dependencies through simulated interventions. For instance, if we consider a sequence $S_{1}$ of two labels $Y_{1},Y_{2}$ with the MB above, we could perform counterfactual interventions by applying $do(X_{1}=0),do(X_{3}=0)$to $S_{1}$. Then we would observe the average change in the likelihood of $Y_{1}$ which, if it is non-zero, would indicate a dependence between $Y_{1}$ and $Y_{2}$. learningcommoncausalvarlabel points out that the assumptions of these inter-label dependencies are already anchored in the Markov Boundaries; we do the same here.
NADEs. Due to the usage of flexible NADEs, we can relax common assumptions regarding data generation processes, such as Poisson Processes or SCMs. Finally, as seen in the Ablations G.1, the effectiveness of OSCAR hinges on the capacity of $\text{Tf}_{x}$ and $\text{Tf}_{y}$ to approximate true conditional probabilities (A4) and provide Oracle CI-test. While assuming Oracle tests are common in the literature [mbb-by-mbb, feature_selection_review] and necessary to recover correct causal structures, this remains a strong assumption. And it is only valid to the extent that the models are perfectly trained. Especially for multi-label classification, performance may degrade in underrepresented regions of the data distribution, as we saw during the MB length comparison.
@@ Appendix I Figures
The following is the implementation of OSCAR in PyTorch [pytorch].